# Synchrotron break frequencies of mildly-to-highly relativistic outflows observed off-axis

### v1

@@ Abstract
keywords:
(transients:) neutron star mergers – (transients:) black hole-neutron star mergers – gravitational waves – radio continuum: transients – relativistic processes
pubyear: 2024pagerange: Synchrotron break frequencies of mildly-to-highly relativistic outflows observed off-axisA
@@ 1 Introduction
As mentioned in Linial & Sari (2019), relativistic electrons undergo energy loss through two distinct mechanisms: (i) synchrotron emission and (ii) up-scattering of photons by inverse-Compton processes. The latter process relies on the presence of low-energy photons. During the initial 10 days, the KN serves as such a photon source, necessitating consideration of inverse-Compton cooling.
@@
Off-axis relativistic jets with an observed afterglow in which a gamma-ray burst is not detected are known as orphan afterglows (Granot et al., 2002). These events are among the promising electromagnetic counterparts to compact objects merger (Metzger & Berger, 2012). Recently, such an afterglow was observed to have a Lorentz factor of $\gamma<10$ (Perley et al., 2024). Additionally, there are indications that tidal disruption events can also generate mildly-to-highly relativistic outflows (Burrows et al., 2011; Saxton et al., 2017; Alexander et al., 2020; Andreoni et al., 2022; Pasham et al., 2023; Rhodes et al., 2023). Providing simple analytic expressions for the break frequencies for these types of outflows is essential for accurately determining their properties.
@@
We derive the synchrotron break frequencies for a uniform volume element of shocked plasma propagating at mildly to highly relativistic velocities, considering both on-axis and off-axis observers. In this description, we assume that the velocity direction of the shocked plasma is perpendicular to the shock front, or equivalently, $v_{\perp}\gg v_{\parallel}$, where $v_{\perp}/v_{\parallel}$ is the lab frame shocked plasma velocity component perpendicular/parallel to the shock front. It applies to various relativistic astrophysical phenomena, including: on-axis emission from a relativistic jet, off-axis emission from a structured jet for as long as the structure propagates radially (Govreen-Segal & Nakar (2023) showed that a power-law jet structure with $dE/d\theta\propto\theta^{\gtrsim-3}$, keeps its initial shape for $\gamma\geq 3$, indicating that the evolution of each angular bin can be approximated by Blandford-McKee evolution (Blandford & McKee, 1976) such that the lateral spreading is negligible), off-axis emission from a relativistic jet as the Lorentz factor reaches $\gamma\sim 1/\theta_{v}$ (where $\theta_{v}$ is the viewing angle) and for a shocked plasma driven by ejecta that propagates into the ISM, such that the shock propagation can be approximated by a section of a spherical outflow (Sadeh, 2024). For the self-absorption frequency, a specific geometrical structure is assumed. We further consider a power-law energy distribution for electrons, given by $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. Here, $\gamma_{e}$ represents the electron Lorentz factor within the plasma’s rest frame, while $p$ is within the range of $2\leq p\leq 2.5$. Simultaneously, fractions $\varepsilon_{e}$ and $\varepsilon_{B}$ correspond to the post-shock internal energy density proportions attributed to non-thermal electrons and magnetic fields, respectively. This phenomenological description, capturing the post-shock plasma conditions, finds support across a diverse spectrum of observations and plasma calculations encompassing both relativistic and non-relativistic shocks (see Blandford & Eichler, 1987; Keshet & Waxman, 2005; Waxman, 2006; Keshet, 2006; Bykov & Treumann, 2011; Sironi et al., 2013; Pohl et al., 2020; Ligorini et al., 2021; Kobzar et al., 2021).
@@ 2.1 Peak frequency
$\nu_{m}=0.7\left(\frac{l_{p}(\gamma-1)}{p-1}\frac{\varepsilon_{e}m_{p}}{m_{e}}\right)^{2}\sqrt{\frac{\gamma-1}{\gamma}}\cdot\frac{q_{e}\sqrt{\varepsilon_{B}nm_{p}}}{m_{e}(1-\beta\cos(\theta_{e}))}.$ (6)
@@ 2.2 Self-absorption frequency
To derive the self-absorption frequency in the observer frame (primed, , quantities are in the shocked plasma frame), we first approximate the absorption coefficient, which is given by (Rybicki & Lightman, 1979)
@@
$f_{a}(p)=(p-1)\left(\frac{p-2}{p-1}\frac{\frac{\varepsilon_{e}m_{p}}{m_{e}}}{1-\left(\gamma_{\text{max}}/\gamma_{\text{min}}\right)^{2-p}}\right)^{p-1}4n\left(\frac{2\pi m_{e}c}{3q_{e}}\right)^{-\frac{p}{2}}\times$ (11)
$\frac{(32\pi\varepsilon_{B}nm_{p}c^{2})^{\frac{p+2}{4}}\sqrt{3}q_{e}^{3}}{8\pi m_{e}^{2}c^{2}}\frac{\sqrt{\pi}}{2}\frac{\Gamma\left(\frac{p+6}{4}\right)}{\Gamma\left(\frac{p+8}{4}\right)}\Gamma\left(\frac{3p+22}{12}\right)\Gamma\left(\frac{3p+2}{12}\right).$
@@
$\tau_{\nu}\approx\alpha_{\nu}\Delta_{\tau}\approx\frac{\beta ct\gamma^{\frac{2-p}{4}}\left(1-\beta\cos(\theta_{e})\right)^{-\frac{p+2}{2}}(\gamma-1)^{\frac{5p-2}{4}}f_{a}(p)\nu^{-\frac{p+4}{2}}}{(12\gamma^{2}\cos(\theta_{e}))(1-\beta\cos(\theta_{e}))}.$ (12)
@@ 2.3 Synchrotron Cooling frequency
$h\nu_{c}^{\text{syn}}=0.7\left(\frac{\left(1-\beta\cos(\theta_{e})\right)m_{e}}{4(\gamma-1)\varepsilon_{B}\left(\frac{4}{3}\right)\sigma_{T}nm_{p}ct}\right)^{2}\sqrt{\frac{\gamma-1}{\gamma}}\cdot\frac{hq_{e}\sqrt{\varepsilon_{B}nm_{p}}}{m_{e}(1-\beta\cos(\theta_{e}))}.$ (15)
@@ 2.4 Inverse Compton cooling
$h\nu_{c}^{\text{IC}}=0.7\left(\frac{\gamma^{3}\beta^{2}3\pi m_{e}c^{4}t}{\sigma_{T}L_{\text{bol}}\left(1-\beta\cos(\theta_{e})\right)}\right)^{2}\sqrt{\frac{\gamma-1}{\gamma}}\cdot\frac{hq_{e}\sqrt{\varepsilon_{B}nm_{p}}}{m_{e}(1-\beta\cos(\theta_{e}))}.$ (19)
@@ 4.2 Previous fits to GW170817
Some of the previous fits to the synchrotron observations of GW170817 provide values in tension with the abovementioned bounds. For example, in many semi-analytical calculations, the best fits found indicate $\varepsilon_{B,-2}n_{-3}=0.1-1$, suggesting that it is essential that the bounds from the spectrum analysis should be incorporated into the fitting pipeline. Furthermore, it is important to notice that the potential jet structure that fits the rising non-thermal light curve of GW170817 is consistent with the lower bounds over the the shocked plasma Lorentz factor, $\gamma>2.6$.
@@ 5 Conclusions
Applying our derived analytic formulae for the case of GW170817 provided bounds over the ISM density, $n$, and the fraction of energy held by magnetic fields $\varepsilon_{B}$ such that $n\cdot\varepsilon_{B}\leq 3\times 10^{-7}\text{cm}^{-3}$. The Lorentz factor of the emitting shocked plasma is bounded by $\gamma>2.6$ at $t\sim 10$days and $\gamma<9.5$ at $t<230$days. Some of the results trying to fit the synchrotron emission data of GW170817 are in tension with these bounds (see §4). For any robust attempt to fit future similar events, it’s essential to take into account the synchrotron break frequencies. This consideration is crucial to ensure the model’s validity, particularly in light of the non-trivial constraints discussed above.

### v2

@@ Abstract
X-ray transient sources(1852)– Radio transient sources(2008)– Gravitational wave sources(677) – Neutron stars(1108)– Relativistic fluid dynamics(1389)
@@ 1 Introduction
As mentioned in Linial & Sari (2019), relativistic electrons undergo energy loss through two distinct mechanisms: (i) synchrotron emission and (ii) up-scattering of photons by inverse-Compton processes. The latter process relies on the presence of low-energy photons. During the initial 10 days, the KN serves as such a photon source, necessitating consideration of inverse-Compton cooling. In general, the low-energy tail of the synchrotron emission also serves as a source of seed photons for synchrotron self-Compton (SSC) processes. However, the potential upper limit for the cooling frequency derived from SSC is significantly higher. This is because the cooling frequency is inversely proportional to the square of the photon energy density, which is dominated by the KN.
@@
Off-axis relativistic jets with an observed afterglow in which a gamma-ray burst is not detected are known as orphan afterglows (Granot et al., 2002). These events are among the promising electromagnetic counterparts to compact objects merger (Metzger & Berger, 2012). Recently, such an afterglow was observed to have a Lorentz factor of $\gamma<10$ (Perley et al., 2024). Additionally, there are indications that tidal disruption events can also generate mildly-to-highly relativistic outflows (Burrows et al., 2011; Saxton et al., 2017; Alexander et al., 2020; Andreoni et al., 2022; Beniamini et al., 2023; Rhodes et al., 2023). Providing simple analytic expressions for the break frequencies for these types of outflows is essential for accurately determining their properties.
@@
We derive the synchrotron break frequencies for a uniform volume element of shocked plasma propagating at mildly to highly relativistic velocities, considering both on-axis and off-axis observers. In this description, we assume that the velocity direction of the shocked plasma is perpendicular to the shock front (the normal), or equivalently, $v_{\perp}\gg v_{\parallel}$, where $v_{\perp}/v_{\parallel}$ is the lab frame shocked plasma velocity component perpendicular/parallel to the shock front. It applies to various relativistic astrophysical phenomena, including: on-axis emission from a relativistic jet, off-axis emission from a structured jet for as long as the structure propagates radially (Govreen-Segal & Nakar (2023) showed that a power-law jet structure with $dE/d\theta\propto\theta^{\gtrsim-3}$, keeps its initial shape for $\gamma\geq 3$, indicating that the evolution of each angular bin can be approximated by Blandford-McKee evolution (Blandford & McKee, 1976) such that the lateral spreading is negligible), off-axis emission from a relativistic jet as the Lorentz factor reaches $\gamma\sim 1/\theta_{v}$ (where $\theta_{v}$ is the viewing angle) and for a shocked plasma driven by ejecta that propagates into the ISM, such that the shock propagation can be approximated by a section of a spherical outflow (Sadeh et al., 2024). For the self-absorption frequency, a specific geometrical structure is assumed. We further consider a power-law energy distribution for electrons, given by $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. Here, $\gamma_{e}$ represents the electron Lorentz factor within the plasma’s rest frame, while $p$ is within the range of $2\leq p\leq 2.5$. Simultaneously, fractions $\varepsilon_{e}$ and $\varepsilon_{B}$ correspond to the post-shock internal energy density proportions attributed to non-thermal electrons and magnetic fields, respectively. This phenomenological description, capturing the post-shock plasma conditions, finds support across a diverse spectrum of observations and plasma calculations encompassing both relativistic and non-relativistic shocks (see Blandford & Eichler, 1987; Keshet & Waxman, 2005; Waxman, 2006; Keshet, 2006; Bykov & Treumann, 2011; Sironi et al., 2013; Pohl et al., 2020; Ligorini et al., 2021; Kobzar et al., 2021).
@@ 2.1 Peak frequency
$\nu_{m}$ $=0.7\left(\frac{l_{p}(\gamma-1)}{p-1}\frac{\varepsilon_{e}m_{p}}{m_{e}}\right)^{2}\times$ (6)
$\sqrt{\frac{\gamma-1}{\gamma}}\cdot\frac{q_{e}\sqrt{\varepsilon_{B}nm_{p}}}{m_{e}(1-\beta\cos(\theta_{e}))}.$
@@ 2.2 Self-absorption frequency
The $1/12$ factor is accurate to $\sim 30\%$ for $\gamma>2$ shocked plasma. It is valid for both a forward-reverse shock structure, and for a relativistic blast wave propagating radially (without an ejecta propagation behind it). To derive the self-absorption frequency in the observer frame (primed, , quantities are in the shocked plasma frame), we first approximate the absorption coefficient, which is given by (Rybicki & Lightman, 1979)
@@
$f_{a}(p)$ $=(p-1)\left(\frac{p-2}{p-1}\frac{\frac{\varepsilon_{e}m_{p}}{m_{e}}}{1-\left(\gamma_{\text{max}}/\gamma_{\text{min}}\right)^{2-p}}\right)^{p-1}\times$ (11)
$4n\left(\frac{2\pi m_{e}c}{3q_{e}}\right)^{-\frac{p}{2}}\frac{(32\pi\varepsilon_{B}nm_{p}c^{2})^{\frac{p+2}{4}}\sqrt{3}q_{e}^{3}}{8\pi m_{e}^{2}c^{2}}\times$
$\frac{\sqrt{\pi}}{2}\frac{\Gamma\left(\frac{p+6}{4}\right)}{\Gamma\left(\frac{p+8}{4}\right)}\Gamma\left(\frac{3p+22}{12}\right)\Gamma\left(\frac{3p+2}{12}\right).$
@@
$\tau_{\nu}$ $\approx\alpha_{\nu}\Delta_{\tau}\approx$ (12)
$\frac{\beta ct\gamma^{\frac{2-p}{4}}\left(1-\beta\cos(\theta_{e})\right)^{-\frac{p+2}{2}}(\gamma-1)^{\frac{5p-2}{4}}f_{a}(p)\nu^{-\frac{p+4}{2}}}{(12\gamma^{2}\cos(\theta_{e}))(1-\beta\cos(\theta_{e}))}.$
@@ 2.3 Synchrotron Cooling frequency
$h\nu_{c}^{\text{syn}}$ $=0.7\left(\frac{\left(1-\beta\cos(\theta_{e})\right)m_{e}}{4(\gamma-1)\varepsilon_{B}\left(\frac{4}{3}\right)\sigma_{T}nm_{p}ct}\right)^{2}\times$ (15)
$\sqrt{\frac{\gamma-1}{\gamma}}\cdot\frac{hq_{e}\sqrt{\varepsilon_{B}nm_{p}}}{m_{e}(1-\beta\cos(\theta_{e}))}.$
@@ 2.4 Inverse Compton cooling
$h\nu_{c}^{\text{IC}}$ $=0.7\left(\frac{\gamma^{3}\beta^{2}3\pi m_{e}c^{4}t}{\sigma_{T}L_{\text{bol}}\left(1-\beta\cos(\theta_{e})\right)}\right)^{2}\times$ (19)
$\sqrt{\frac{\gamma-1}{\gamma}}\cdot\frac{hq_{e}\sqrt{\varepsilon_{B}nm_{p}}}{m_{e}(1-\beta\cos(\theta_{e}))}.$
@@ 4.2 Previous fits to GW170817
Some of the previous fits to the synchrotron observations of GW170817 provide values in tension with the abovementioned bounds. For example, in many semi-analytical calculations, the best fits found indicate $\varepsilon_{B,-2}n_{-3}=0.1-1$, suggesting that it is essential that the bounds from the spectrum analysis should be incorporated into the fitting pipeline. Furthermore, it is important to notice that the potential jet structure that fits the rising non-thermal light curve of GW170817 is consistent with the lower bounds over the shocked plasma Lorentz factor, $\gamma>2.6$. These bounds are consistent with the suggested alternative in Sadeh & Waxman (2024), a conical, radially-stratified outflow.
@@ 5 Conclusions
Applying our derived analytic formulae for the case of GW170817 provided bounds over the ISM density, $n$, and the fraction of energy held by magnetic fields $\varepsilon_{B}$ such that $n\cdot\varepsilon_{B}\leq 3\times 10^{-7}\text{cm}^{-3}$. The Lorentz factor of the emitting shocked plasma is bounded by $\gamma>2.6$ at $t\sim 10$ days and $\gamma<9.5$ at $t<230$ days. Some of the results trying to fit the synchrotron emission data of GW170817 are in tension with these bounds (see §4), while it is consistent with the solution suggested in Sadeh & Waxman (2024). For any robust attempt to fit future similar events, it’s essential to take into account the synchrotron break frequencies. This consideration is crucial to ensure the model’s validity, particularly in light of the non-trivial constraints discussed above.
The effect of the thermal population of electrons has been explored in Margalit & Quataert (2021, 2024), suggesting it could have a significant impact the emission spectrum at frequencies of $\nu\gtrsim 1$GHz in cases of mildly relativistic shocks propagating in ISM densities of $\gg 1$cm-3 with $\varepsilon_{e}\ll 0.1$. Such conditions are primarily relevant to phenomena like "Fast Blue Optical Transients" (Ho et al., 2022). For instance, the late ($\lesssim 1000$ days) non-thermal emission following GW170817 is consistent with emission from a power-law distribution of electrons due to the relatively low ISM density, $n\ll 1$cm-3, despite the shocked plasma being mildly relativistic. Nevertheless, the thermal population does not influence the derivations obtained in this paper which are valid for the power-law population of electrons. Additionally, in this paper, we considered a constant ISM density, $n$, which may not apply to cases involving substantial mass loss via winds from the progenitor, which is beyond the scope of this paper. Such constant mass-loss wind is expected to produce $n\propto r^{-2}$ profiles altering some of the results discussed in this work. While the peak frequency and inverse Compton cooling frequency would remain unaffected, the self-absorption and synchrotron cooling frequencies would increase and decrease, respectively, due to the higher densities and, correspondingly, higher magnetic fields encountered during the expansion.